undeadknight26 (undeadknight26):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Medal?

OpenStudy (anonymous):

f(5) = 2 + 5 + f(4) F(4) is twice f(2) so I say it is 10 f(5 ) = 2 + 5+ 10 = 17

undeadknight26 (undeadknight26):

Thank you. can you help me with one more?

OpenStudy (anonymous):

sure

undeadknight26 (undeadknight26):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

let me go back to the first one f(4) = 25 so f(5) = 2 + 5 + 25 = 32

undeadknight26 (undeadknight26):

So its 32?

OpenStudy (anonymous):

0 1 1 2 3

OpenStudy (anonymous):

f(1 ) = 0 given f(2) = 1 given f(3) = f(2) + f(1) = 1 + 0 = 1 where n = 3, just substitute n and solve f(4) = f(3) + f(2) = 1 + 1 = 2 , where n = 4 f(5) = f(4) + f(3) = 2 + 1 = 3 , where n = 5

undeadknight26 (undeadknight26):

thank you

OpenStudy (anonymous):

you are welcome